A fighter jet is catapulted off an aircraft carrier from rest to 75 m/s. If the aircraft carrier deck is 100 m long,

what is the acceleration of the jet?


1
Expert's answer
2021-09-13T11:01:04-0400

According to the kinematic equation, the distance travelled by the fighter jet is given as follows (see https://www.khanacademy.org/science/physics/one-dimensional-motion/kinematic-formulas/a/what-are-the-kinematic-formulas):


"\Delta x = \dfrac{v_f^2 - v_i^2}{2a}"

where "v_f = 75m\/s, v_i = 0m\/s" are final and initial speeds of the jet respectively, "a" is its acceleration. Assuming "\Delta x = 100m", obtain:


"a = \dfrac{v_f^2 - v_i^2}{2\Delta x} \\
a = \dfrac{75^2 -0^2}{2\cdot 100} = 28.125 m/s^2"

Answer. "28.125 m\/s^2".